OpenStudy (anonymous):

WHAT ARE THE FACTORS OF THE FOLLOWING TRINOMIAL? F^2 - 14F -49

OpenStudy (anonymous):

I'd probably use the quadratic formula or complete the square to find the factors

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

is it - 49 or + 49 at the end??/

OpenStudy (anonymous):

+ 49

OpenStudy (anonymous):

F^2 - 14F + 49 this can be written as (F)^2 - 2(F)(7) + (7)^2 Since a^2 - 2ab + b^2 = (a - b)^2 Hence (F)^2 - 2(F)(7) + (7)^2 = (F - 7)^2 => F^2 - 14F + 49 = (F - 7)^2 = (F - 7)(F-7)